OpenStudy (anonymous):

Consider the equation: 2cos^2(theta) - 2cos(theta) -1=0. I've used the quadratic formula to find the degree solutions but it appears that I have done it wrong. I'm trying to find the degree solutions. I know the answer is __?__degrees + 360 degree K and there are two answers but I am having a hard time arriving at the answer. Can someone please help me? I'd appreciate it.

OpenStudy (campbell_st):

use substitution let \[u = \cos(\theta)\] then 2u^2 - 2u - 1 = 0 using the general quadratic formula \[u=\frac{2\pm \sqrt{(-2)^2 - 4 \times2\times(-1)}}{2\times2}\] \[u = \frac{2\pm2\sqrt{3}}{4}\] \[u = \frac{1\pm \sqrt{3}}{2}\] resubstituting \[cos(\theta) = \frac{1\pm \sqrt{3}}{2}\]

OpenStudy (anonymous):

I get the decimal answers for 1+- Sq.rt3/2 and plug it into the original equation?

OpenStudy (shayaan_mustafa):

Hi jennifer473! campbell_st has done it for you. still confusion?

OpenStudy (anonymous):

I understand the quadratic formula. from the answer I get using the quadratic formula (which is a decimal), do I plug it into the original formula? (2cos^2(theta) - 2cos(theta)-1=0)

OpenStudy (shayaan_mustafa):

no. In question you are asking for solution. so campbell_st done it for you. If you plug your solution in the question then it will be your verification.

OpenStudy (anonymous):

So then how do I use the answer formed when using the quadratic equation to get the degrees?

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

Yes please :)

OpenStudy (anonymous):

from what I'm gathering you want the angle theta from campbell_st 's solution... notice there are two solutions... for the first one, cos(theta) = (1+sqrt3)/2. this does not have a solution because that right side is bigger than 1. so we can throw this one away.

OpenStudy (anonymous):

for the other one, cos(theta) = (1-sqrt3)/2, the right side is between -1 and 1 so this has a solution. to find theta, just take the cosine inverse of that right side: